OpenStudy (anonymous):

Assuming A≠0, evaluate the limit lim 3sin^2(At) / cos(At) - 1 t-->0

OpenStudy (anonymous):

Start with: \[\sin^2(\theta) + \cos^2(\theta) = 1\] Solve that for sin^2(theta), and substitute

OpenStudy (anonymous):

use 1 - cos At = 2sin^2(At/2 ) also sin^2(At) = 4sin^2(At/2)*cos^2(At/2

OpenStudy (anonymous):

lim 3sin^2(At) / cos(At) - 1 t-->0 = lim - 3 *4sin^2(At/2)*cos^2(At/2/ 2sin^2(At/2 ) t-->0 now you can do further

OpenStudy (anonymous):

Alternate way: \[3\frac{\sin^2(At)}{\cos(At) - 1} = 3\frac{\cos^2(At) - 1}{\cos(At) - 1} = 3\frac{(\cos(At)-1)(\cos(At)+1)}{\cos(At)-1}\]

OpenStudy (anonymous):

Oh, small mistake. Bleh. Should be: \[3\frac{1-\cos^2(At)}{\cos(At)-1}\] for the middle part, and: \[3\frac{(1-\cos(At))(1+\cos(At))}{(\cos(At) - 1)} = -3\frac{(\cos(At) - 1)(1+\cos(At))}{(\cos(At) - 1)}\]

OpenStudy (anonymous):

thanks :)